OpenStudy (anonymous):

Factor the trinomial below. Write each factor as a polynomial in descending order. 4x2 + 19x - 5 Please :)

Directrix (directrix):

Look for two numbers that multiply to -20 and add to 19. -20 and 1 will do. 4x2 + 19x - 5 = 4x^2 + 1x - 20x -5 = x(4x-1) - 5(4x +1) = Factor out the common factor of (4x -1) = (4x -1) ( x - ?) @rayado21